# A Survey on Unlearning in Large Language Models

### v1

@@ 4.2.2. Model Utility
Beyond investigating model memorization, various methodologies are employed to assess the general utility of unlearned models. Some directly and efficiently computable indicators quantify specific aspects of model performance, which is referred to as utility metrics. Among these, Perplexity is one of the most frequently used measures; a lower perplexity signifies better fluency (Liu et al., 2021), higher model confidence (Wei et al., 2025b), and improved meaningfulness of the generated content (Liu et al., 2024a). In addition to perplexity, numerous studies focus on lexical diversity, proposing metrics such as the mean number of distinct n-grams (Liu et al., 2021), the unique token ratio (Liu et al., 2024a; Yao et al., 2024b), and token entropy (Yuan et al., 2025b). As noted by Yuan et al. (2025b), reduced vocabulary diversity often correlates with token repetition in model outputs, indicating poorer readability and weaker overall utility. Also, established linguistic indices, including Brunet’s Index (Brunet et al., 1978) and Honore’s Statistic (Honoré et al., 1979), have been applied to assess lexical richness in unlearning contexts (Xu et al., 2025).
@@
Meanwhile, comprehensive general benchmarks are commonly utilized to evaluate the overall performance of unlearning methods (Jin et al., 2024; Li et al., 2024). Table 6 summarizes frequently adopted benchmarks and their usage statistics across different unlearning studies. Integrated evaluation frameworks and toolkits, such as Language Model Evaluation Harness (Gao et al., 2021), facilitate the systematic application of these benchmarks for LLM unlearning evaluation (Jia et al., 2025; Ashuach et al., 2025; Wang et al., 2025g; Tian et al., 2024).
     Name Description Used by
     ARC (Clark et al., 2018) Reasoning (Eldan and Russinovich, 2024; Yao et al., 2024a; Liu et al., 2025b; Yuan et al., 2025b; Chowdhury et al., 2025; Thaker et al., 2024; Huang et al., 2025; Ji et al., 2024; Wang et al., 2025e, g; Tian et al., 2024; Gu et al., 2024; Veldanda et al., 2024; Choi et al., 2025)
     GSM8K (Cobbe et al., 2021) Math (Yao et al., 2024a; Ashuach et al., 2025; Yuan et al., 2025b; Hu et al., 2024; Wang et al., 2025g)
     HellaSwag (Zellers et al., 2019) Commonsense (Eldan and Russinovich, 2024; Chowdhury et al., 2025; Thaker et al., 2024; Huang et al., 2025; Ji et al., 2024; Wang et al., 2025g; Choi et al., 2025; Russinovich and Salem, 2025)
     MMLU (Hendrycks et al., 2021) Universal knowledge (Jin et al., 2024; Li et al., 2024; Yao et al., 2024a; Ashuach et al., 2025; Ramakrishna et al., 2025; Lang et al., 2025; Che et al., 2025; Doshi and Stickland, 2024; Yuan et al., 2025b; Farrell et al., 2024; Hu et al., 2024; Chowdhury et al., 2025; Ren et al., 2025; Thaker et al., 2024; Liu et al., 2024a; Sanyal and Mandal, 2025; Wang et al., 2025e; Dou et al., 2025; Hu et al., 2025; Guo et al., 2025; Dang et al., 2025; Tian et al., 2024; Choi et al., 2025; Zhang et al., 2025e; Łucki et al., 2025; Tamirisa et al., 2024; Russinovich and Salem, 2025; Wei et al., 2025a)
     MT-Bench (Zheng et al., 2023) Multi-turn (Li et al., 2024; Lang et al., 2025; Che et al., 2025; Doshi and Stickland, 2024; Thaker et al., 2024; Dou et al., 2025; Russinovich and Salem, 2025; Wei et al., 2025a)
@@
     OpenBookQA (Mihaylov et al., 2018) Understanding & application (Eldan and Russinovich, 2024; Gao et al., 2025; Chowdhury et al., 2025; Thaker et al., 2024; Huang et al., 2025; Ji et al., 2024; Choi et al., 2025)
     PIQA (Bisk et al., 2020) Commonsense (Eldan and Russinovich, 2024; Liu et al., 2025b; Chowdhury et al., 2025; Ji et al., 2024; Gu et al., 2024; Choi et al., 2025)
@@
     TruthfulQA (Lin et al., 2022) Mimic human Falsehoods (Jin et al., 2024; Yuan et al., 2025b; Hu et al., 2024; Chowdhury et al., 2025; Tian et al., 2024; Zhang et al., 2025e; Yuan et al., 2025a; Russinovich and Salem, 2025)
     WinoGrande (Sakaguchi et al., 2021) Commonsense (Eldan and Russinovich, 2024; Liu et al., 2025b; Chowdhury et al., 2025; Huang et al., 2025; Choi et al., 2025; Russinovich and Salem, 2025)
@@
Table 6. Overview of general benchmarks widely used in unlearning evaluation.
@@ 4.2.3. Unlearning Robustness
Empirical studies indicate that many machine methods merely suppress the surface-level expression of specific knowledge, leaving the underlying representations vulnerable to various adversarial attacks (Lynch et al., 2024; Hong et al., 2024; Łucki et al., 2025). To systematically assess robustness, a range of adversarial techniques from the security domain have been integrated into the evaluation of LLM unlearning (Lynch et al., 2024; Hong et al., 2024; Li et al., 2024; Che et al., 2025; Sanyal and Mandal, 2025; Jin et al., 2024; Shi et al., 2025; Ramakrishna et al., 2025; Jia et al., 2024), which we collectively refer to as attack techniques. Commonly adopted methods include: (1) Attack on the input, such as crafted jailbreak prompts (Shen et al., 2024a), in-context learning adversarial attacks (Wei et al., 2023), GCG (Zou et al., 2023), AutoPrompt (Shin et al., 2020), BEAST (Sadasivan et al., 2024), PAIR (Chao et al., 2025), persona modulation (Shah et al., 2023), JailbreakHub (Shen et al., 2024b), many-shot jailbreaking (Anil et al., 2025). (2) Attack on the hidden layer, such as probing techniques (Belinkov, 2022; Alain and Bengio, 2016), soft-prompt-based threats (Schwinn et al., 2024; Che et al., 2025), AnonAct (Seyitoğlu et al., 2024), Logit Lens (nostalgebraist, 2020). (3) Membership inference attack (MIA), which is a privacy attack that determines whether specific data samples are part of a model’s training set (Shokri et al., 2017), including LOSS (Yeom et al., 2018), Zlib Entropy (Carlini et al., 2021), Min-K% Prob (Shi et al., 2024) and Min-K%++ Prob (Zhang et al., 2025c). Further heuristic attacks have also been proposed for specific unlearning scenarios (Ashuach et al., 2025; Shen et al., 2025; Du et al., 2025).
In response to the characteristics of LLM unlearning tasks, a relatively unique robustness evaluation method, relearning (Lynch et al., 2024; Joshi et al., 2024), is also frequently used. Relearning evaluates an unlearned model by exposing it to a limited subset of the unlearned data. In in-context relearning, knowledge related to the unlearn set, such as book summaries or relevant background information, are included in the prompts when evaluating the unlearned LLM. When relearning by fine-tuning, the model is full-parameter or LoRA fine-tuned on a small portion of the unlearn set or a related set. Empirical studies consistently demonstrate that relearning can substantially degrade unlearning quality, causing the model to rapidly recapitulate a significant portion of unlearned knowledge from sparse cues (Lynch et al., 2024; Lermen and Rogers-Smith, 2024; Zhan et al., 2024; Guo et al., 2025; Joshi et al., 2024), or begin to systematically avoid generating content related to the unlearning target, even when contextually prompted (Joshi et al., 2024). More critically, Doshi and Stickland (2024) reveal that fine-tuning on an entirely benign dataset can also reverse the unlearning effects, restoring model performance to a level comparable to its state before unlearning.
@@ 4.2.4. Unlearning efficiency
While the majority of existing studies concentrate on the efficacy of unlearning, the resource overhead of deploying such algorithms under real-world constraints remains a crucial consideration, including both memory occupation (Bannihatti Kumar et al., 2023; Jia et al., 2024) and computational time consumption. For time cost, a straightforward approach is to directly measure the algorithm’s runtime during experiments (Chen and Yang, 2023; Chowdhury et al., 2025; Bannihatti Kumar et al., 2023; Jia et al., 2024). Since computational speed is closely tied to GPU performance, some studies convert raw runtime into GPU hours (i.e., number of GPUs × training hours) to facilitate comparison (Ren et al., 2025; Ji et al., 2024). Nonetheless, fair cross-study comparisons remain challenging due to variations in experimental environments. The most reliable method is to execute all algorithms under controlled conditions, though this is often resource-intensive. Alternatively, several works estimate time consumption theoretically, using metrics such as floating-point operations (Cha et al., 2024) or gradient computation budgets (Wei et al., 2025b). However, discrepancies between theoretical estimates and actual runtime may arise due to differences in implementation and hardware optimization.
@@ 5.1.1. Definition and Evaluation of Unlearning
(1) How should memorization be defined and detected? Most studies assess memorization based on model output in specific tasks, yet disagree on the criteria for judging these outputs. For the content related to the unlearn set, some works argue that the model should simply avoid generating such content (Eldan and Russinovich, 2024), while others require it to explicitly respond with “I don’t know” (Shen et al., 2025). Another line of research proposes that the unlearned model should produce outputs similar to those of a hypothetical retrained model, such as giving a specific incorrect answer (Rezaei et al., 2024). When direct output is insufficient, adversarial methods are sometimes employed to expose memorization. However, such approaches face inherent limitations: Overly weak attacks may fail to detect memorization, whereas overly strong ones can force the model to generate arbitrary content, casting doubt on their reliability as auditing tools (Chen et al., 2025). Moreover, certain attack methods are considered ill-suited to the LLM context, such as MIA, which typically requires training numerous shadow models, thus being both data-prohibitive and computationally intractable for LLMs (Liu et al., 2024a).
(2) What should constitute the unlearn set? For synthetic datasets such as TOFU (Maini et al., 2024), this question is relatively straightforward. However, in real-world scenarios, data interconnectivity complicate the identification of appropriate unlearning targets. Tian et al. (2024) adopt a legal perspective to determine which copyrighted or private data should be unlearned, while Wu et al. (2025) construct relationship graphs to identify necessary data for removal. Despite their merits, these methods remain reliant on manual, domain-specific analysis, and lack generalizability.
@@ 5.1.2. Effect of Unlearning
Effects across languages. Some studies conduct evaluations with prompts translated into languages other than English, finding that monolingual unlearning is fundamentally insufficient for multilingual LLMs (Lynch et al., 2024; Jin et al., 2024). Furthermore, more languages that systematically divide into high- and low-resource are used in evaluations (Choi et al., 2024a; Lu and Koehn, 2025), revealing the fact that unlearning in one language does not necessarily transfer to others and could even inadvertently reinforce harmful content across languages. Together, these findings underscore a critical consensus: effective and secure unlearning necessitates multilingual joint unlearning strategies that are designed to address model behavior holistically in all languages.
Effects across data. From the perspective of data distribution, Baluta et al. (2024) demonstrate that out-of-distribution (OOD) data require more gradient ascent but offer a better unlearning quality, whereas in-distribution data allow faster unlearning but severely compromise model utility, illustrating a fundamental trade-off between unlearning efficiency and model preservation. Considering the logical connectivity of the data, Choi et al. (2024b) identify that current unlearning methods struggle with multi-hop knowledge, where unlearning one intermediate fact in a chain often fails to remove the entire logical sequence. Furthermore, some studies investigate the impact on adjacent data after performing unlearning on selected data, identifying phenomena called “transfer unlearning” (Lu et al., 2024), “ripple effect” (Zhang et al., 2025d) and “onion effect” (Borkar, 2023). These effects highlight the intricate and unpredictable consequences of unlearning, emphasizing the need for careful monitoring to ensure that unlearning achieves its intended goals without introducing new risks.
@@ 5.1.3. Unlearning in Reality
A significant challenge lies in the scaling gap between experimental settings and real-world conditions. Current unlearning experiments are largely limited to models with fewer than 10 billion parameters and unlearning sets under 1 billion instances, raising concerns about the applicability of these methods to larger models and datasets. Shi et al. (2025) analyze how evaluation metrics evolve as the size of the unlearn set increases, providing insight into scalability. On the other hand, in practical deployments, large models are often compressed, such as using quantization for efficiency. Notably, Zhang et al. (2025e) demonstrate that quantizing unlearned models can inadvertently reactivate unlearned knowledge, highlighting a key scalability challenge.
In commercial applications, unlearning requests typically arrive sequentially, requiring models to continuously unlearning while maintaining performance (Rezaei et al., 2024; Veldanda et al., 2024). To assess long-term viability, Shi et al. (2025) collect model checkpoints after processing each sequential request and track evaluation metrics over time. This approach helps quantify the cumulative impact of repeated unlearning and the model’s ability to sustain utility. Unfortunately, current unlearning methods are not yet ready to handle sequential unlearning.

### v2

@@ 4.2.2. Model Utility
Beyond investigating model memorization, various methodologies are employed to assess the general utility of unlearned models. Some directly and efficiently computable indicators quantify specific aspects of model performance, which is referred to as utility metrics. Among these, perplexity is one of the most frequently used measures; a lower perplexity signifies better fluency (liu2021dexperts), higher model confidence (wei2024underestimated), and improved meaningfulness of the generated content (Liu2024Large). In addition to perplexity, numerous studies focus on lexical diversity, proposing metrics such as the mean number or proportion of distinct n-grams (liu2021dexperts; Sinha2024UnStar; Liu2024Revisiting), the unique token ratio (Liu2024Large; Yao2024Large), and token entropy (yuan2024closer; Tang2024Learn). As noted by yuan2024closer, reduced vocabulary diversity often correlates with token repetition in model outputs, indicating poorer readability and weaker overall utility. Also, established linguistic indices, including Brunet’s Index (brunet1978vocabulaire) and Honore’s Statistic (honore1979some), have been applied to assess lexical richness in unlearning contexts (xu2025relearn).
@@
Meanwhile, comprehensive general benchmarks are commonly utilized to evaluate the overall performance of unlearning methods (Jin2024RWKU; Li2024WMDP). Table 5 summarizes frequently adopted benchmarks and their usage statistics across different unlearning studies. Integrated evaluation frameworks and toolkits, such as Language Model Evaluation Harness (gao2021framework), facilitate the systematic application of these benchmarks for LLM unlearning evaluation (jia2024wagle; Ashuach2024REVS; Wang2024Large; Tian2024Forget; Kadhe2024Split).
    Class Name Used by
    Reasoning ARC (clark2018think) (Eldan2023Whos; yao2024machine; Liu2024Learning; yuan2024closer; Chowdhury2024Scalable; thaker2024guardrail; Huang2024Offset; Ji2024Reversing; wang2024when; Wang2024Large; Tian2024Forget; Gu2024MEOW; veldanda2024llm; choi2025opt-out; chen2024wpn; Tang2024Learn; jang2022knowledge; Lee2024Protecting; Wang2024RKLD; barbulescu2024each; Lu2024Eraser; Gu2024SecondOrder; dong2024undial; Wang2024Selective; kassem2023preserving)
    Math GSM8K (cobbe2021training) (yao2024machine; Ashuach2024REVS; yuan2024closer; hu2024separate; Wang2024Large; Lizzo2024UNLEARN)
MathQA (amini2019mathqa) (Gu2024SecondOrder; Tang2024Learn; jang2022knowledge; Lee2024Protecting; Wang2024RKLD; barbulescu2024each; chen2024wpn; Wang2024Selective; kassem2023preserving)
    Commonsense HellaSwag (zellers2019hellaswag) (Eldan2023Whos; Chowdhury2024Scalable; thaker2024guardrail; Huang2024Offset; Ji2024Reversing; Wang2024Large; choi2025opt-out; russinovich2025obliviate; chen2024wpn; Tang2024Learn; jang2022knowledge; Lee2024Protecting; Wang2024RKLD; barbulescu2024each; Lu2024Eraser; dong2024undial; Wang2024Selective; kassem2023preserving)
PIQA (bisk2020piqa) (Eldan2023Whos; Liu2024Learning; Chowdhury2024Scalable; Ji2024Reversing; Gu2024MEOW; choi2025opt-out; chen2024wpn; Tang2024Learn; jang2022knowledge; Lee2024Protecting; Wang2024RKLD; barbulescu2024each; Gu2024SecondOrder; dong2024undial; Wang2024Selective; kassem2023preserving)
WinoGrande (sakaguchi2021winogrande) (Eldan2023Whos; Liu2024Learning; Chowdhury2024Scalable; Huang2024Offset; choi2025opt-out; russinovich2025obliviate; chen2024wpn; Tang2024Learn; jang2022knowledge; Lee2024Protecting; Wang2024RKLD; barbulescu2024each; dong2024undial; Wang2024Selective; kassem2023preserving)
    Comprehension Lambada (paperno2016lambada) (Liu2024Learning; choi2025opt-out; chen2024wpn; Gu2024SecondOrder; Tang2024Learn; jang2022knowledge; Lee2024Protecting; barbulescu2024each; kassem2023preserving)
@@
OpenBookQA (mihaylov2018suit) (Eldan2023Whos; gao2024large; Chowdhury2024Scalable; thaker2024guardrail; Huang2024Offset; Ji2024Reversing; choi2025opt-out)
     Universal knowledge MMLU (hendrycks2020measuring) (Jin2024RWKU; Li2024WMDP; yao2024machine; Ashuach2024REVS; ramakrishna2025lume; lang2025beyond; che2025model; doshi2024does; yuan2024closer; farrell2024applying; hu2024separate; Chowdhury2024Scalable; ren2025general; thaker2024guardrail; Liu2024Large; sanyal2025agents; wang2024when; Dou2024Avoiding; hu2025falcon; guo2024mechanistic; huu2024effects; Tian2024Forget; choi2025opt-out; zhang2025catastrophic; lucki2024adversarial; Tamirisa2024Robust; russinovich2025obliviate; Wei2024Evaluating; Lizzo2024UNLEARN; Kadhe2024Split; deeb2025unlearning)
    Multi-turn MT-Bench (zheng2023judging) (Li2024WMDP; lang2025beyond; che2025model; doshi2024does; thaker2024guardrail; Dou2024Avoiding; russinovich2025obliviate; Wei2024Evaluating)
@@
     Mimic human Falsehoods TruthfulQA (lin2022truthfulqa) (Jin2024RWKU; yuan2024closer; hu2024separate; Chowdhury2024Scalable; Tian2024Forget; zhang2025catastrophic; Yuan2024Robust; russinovich2025obliviate; chen2024machine; Liu2024Safer; Yao2024Large)
@@
Table 5. Overview of general benchmarks widely used in unlearning evaluation. We broadly divide these benchmarks into seven classes and count the use of each benchmark.
@@ 4.2.3. Unlearning Robustness
Empirical studies indicate that many machine methods merely suppress the surface-level expression of specific knowledge, leaving the underlying representations vulnerable to various adversarial attacks (Lynch2024Eight; Hong2024Intrinsic; lucki2024adversarial). To systematically assess robustness, a range of adversarial techniques from the security domain have been integrated into the evaluation of LLM unlearning (Jin2024RWKU; Shi2024MUSE; ramakrishna2025lume; Jia2024SOUL), which we collectively refer to as attack techniques. Commonly adopted methods include the following. (1) Attack on the input, such as crafted jailbreak prompts (shen2024do), AutoPrompt (shin2020autoprompt), Greedy Coordinate Gradient (GCG) (zou2023universal), Prompt Automatic Iterative Refinement (PAIR) (chao2023jailbreaking). (2) Attack on the hidden layer, such as probing techniques (belinkov2022probing; alain2016understanding), Logit Lens (nostalgebraist2020interpreting), soft-prompt-based threats (schwinn2024soft; che2025model), AnonAct (seyitouglu2024extracting). Further heuristic attacks have also been proposed for specific unlearning scenarios (Ashuach2024REVS; shen2025lunar; Du2024Textual).
In response to the characteristics of LLM unlearning tasks, a relatively unique robustness evaluation method, relearning (Lynch2024Eight; joshi2024towards), is also frequently used. Relearning evaluates an unlearned model by exposing it to a limited subset of the unlearned data. In in-context relearning, knowledge related to the unlearn set, such as book summaries or relevant background information, are included in the prompts when evaluating the unlearned LLM. When relearning by fine-tuning, the model is full-parameter or LoRA fine-tuned on a small portion of the unlearn set or a related set. Empirical studies consistently demonstrate that relearning can substantially degrade unlearning quality, causing the model to rapidly recapitulate a significant portion of unlearned knowledge from sparse cues (Lynch2024Eight; lermen2023lora; zhan2023removing; guo2024mechanistic; joshi2024towards; fan2025simplicity), or begin to systematically avoid generating content related to the unlearning target, even when contextually prompted (joshi2024towards). More critically, fine-tuning on data with content or distribution similar to the unlearn set can also reverse the unlearning effects, restoring model performance to a level comparable to its state before unlearning (doshi2024does; deeb2025unlearning).
@@ 4.2.4. Unlearning efficiency
While the majority of existing studies concentrate on the efficacy of unlearning, the resource overhead of deploying such algorithms under real-world constraints remains a crucial consideration, including both memory occupation (BannihattiKumar2023Privacy; Jia2024SOUL) and computational time consumption. For time cost, a straightforward approach is to directly measure the algorithm’s runtime during experiments (Chen2023Unlearn; Chowdhury2024Scalable; BannihattiKumar2023Privacy; Jia2024SOUL). Since computational speed is closely tied to GPU performance, some studies convert raw runtime into GPU hours (i.e., number of GPUs × training hours) to facilitate comparison (ren2025general; Ji2024Reversing). Nonetheless, fair cross-study comparisons remain challenging due to variations in experimental environments. The most reliable method is to execute all algorithms under controlled conditions, though this is often resource-intensive. Alternatively, several works estimate time consumption theoretically, using metrics such as floating-point operations (cha2024robust) or gradient computation budgets (wei2024underestimated). However, discrepancies between theoretical estimates and actual runtime may arise due to differences in implementation and hardware optimization.
@@ 5.1.1. Definition and Evaluation of Unlearning
(1) How should memorization be defined and detected? Most studies assess memorization based on model output in specific tasks, yet disagree on the criteria for judging these outputs. For the content related to the unlearn set, some works argue that the model should simply avoid generating such content (Eldan2023Whos), while others require it to explicitly respond with “I don’t know” (shen2025lunar). Another line of research proposes that the unlearned model should produce outputs similar to those of a hypothetical retrained model, such as giving a specific incorrect answer (rezaei2024restor). When direct output is insufficient, adversarial methods are sometimes employed to expose memorization. However, such approaches face inherent limitations: Overly weak attacks may fail to detect memorization, whereas overly strong ones can force the model to generate arbitrary content, casting doubt on their reliability as auditing tools (chen2025soft).
(2) What should constitute the unlearn set? For synthetic datasets such as TOFU (Maini2024TOFU), this question is relatively straightforward. However, in real-world scenarios, data interconnectivity complicate the identification of appropriate unlearning targets. Tian2024Forget adopt a legal perspective to determine which copyrighted or private data should be unlearned, while wu2024evaluating construct relationship graphs to identify necessary data for removal. Despite their merits, these methods remain reliant on manual, domain-specific analysis, and lack generalizability.
@@ 5.1.2. Effect of Unlearning
Effects across languages. Some studies conduct evaluations with prompts translated into languages other than English, finding that monolingual unlearning is fundamentally insufficient for multilingual LLMs (Lynch2024Eight; Jin2024RWKU). Furthermore, more languages that systematically divide into high- and low-resource are used in evaluations (choi2024cross; lu2025learn), revealing the fact that unlearning in one language does not necessarily transfer to others and could even inadvertently reinforce harmful content across languages. Together, these findings underscore a critical consensus: effective and secure unlearning necessitates multilingual joint unlearning strategies to address model behavior holistically in all languages.
Effects across data. From the perspective of data distribution, baluta2024unlearning demonstrate that out-of-distribution (OOD) data require more gradient ascent but offer a better unlearning quality, whereas in-distribution data allow faster unlearning but severely compromise model utility, illustrating a fundamental trade-off between unlearning efficiency and model preservation. Considering the logical connectivity of the data, choi2024breaking identify that current unlearning methods struggle with multi-hop knowledge, where unlearning one intermediate fact in a chain often fails to remove the entire logical sequence. Furthermore, some studies investigate the impact on adjacent data after performing unlearning on selected data, identifying phenomena called “transfer unlearning” (lu2024transfer), “ripple effect” (zhang2025theft) and “onion effect” (borkar2023learn). These effects highlight the intricate and unpredictable consequences of unlearning, emphasizing the need for careful monitoring to ensure that unlearning achieves its intended goals without introducing new risks.
@@ 5.1.3. Unlearning in Reality
A significant challenge lies in the scaling gap between experimental settings and real-world conditions. Current unlearning experiments are largely limited to models with fewer than 10 billion parameters and unlearning sets under 1 billion instances, raising concerns about the applicability of these methods to larger models and datasets. Shi2024MUSE analyze how evaluation metrics evolve as the size of the unlearn set increases, providing insight into scalability. On the other hand, in practical deployments, large models are often compressed, such as using quantization for efficiency. Notably, zhang2025catastrophic demonstrate that quantizing unlearned models can inadvertently reactivate unlearned knowledge, highlighting a key scalability challenge.
In commercial applications, unlearning requests typically arrive sequentially, requiring models to continuously unlearning while maintaining performance (rezaei2024restor; veldanda2024llm). To assess long-term viability, Shi2024MUSE collect model checkpoints after processing each sequential request and track evaluation metrics over time. This approach helps quantify the cumulative impact of repeated unlearning and the model’s ability to sustain utility. Unfortunately, current unlearning methods are not yet ready to handle sequential unlearning.